### v1

@@ Abstract
Accurately reconstructing a 3D scene including explicit geometry information is both attractive and challenging. Geometry reconstruction can benefit from incorporating differentiable appearance models, such as Neural Radiance Fields and 3D Gaussian Splatting (3DGS). In this work, we propose a learnable scene model that incorporates 3DGS with an explicit geometry representation, namely a mesh. Our model learns the mesh and appearance in an end-to-end manner, where we bind 3D Gaussians to the mesh faces and perform differentiable rendering of 3DGS to obtain photometric supervision. The model creates an effective information pathway to supervise the learning of the scene, including the mesh. Experimental results demonstrate that the learned scene model not only achieves state-of-the-art rendering quality but also supports manipulation using the explicit mesh. In addition, our model has a unique advantage in adapting to scene updates, thanks to the end-to-end learning of both mesh and appearance.
Keywords:
@@
3D reconstruction Gaussian splatting Hybrid representation Mesh-guided manipulation
1 Introduction
@@
Meshes carrying explicit geometric information are often the preferred choice in numerous geometry processing scenarios. When building a 3D scene model that includes a mesh, gradient-based optimization from visual observation is challenging due to the complicated pipeline required to produce the synthetic views. Recently, neural surface reconstruction methods [27, 31] represent scene geometry as an implicit signed distance function (SDF) and establish a connection with Neural Radiance Fields (NeRFs) [17], enabling effective image-based supervision. However, the geometry is learned indirectly from the radiance field and faces the challenge of expensive ray-marching in the volumetric rendering of NeRFs.
@@
3D Gaussian Splatting (3DGS) [14] has recently gained popularity due to its ability to render photorealistic images significantly faster than NeRFs. Despite the advantages in rendering quality and speed, learning a mesh using 3DGS remains problematic due to the misalignment between the Gaussian centers and the shape surface. As illustrated in Fig. 1, the Gaussians form an appearance model that may not accurately represent the geometry required by a mesh representation. Recently, some methods have incorporated regularization into the 3DGS learning process. NeuSG [3] refines Gaussians using surface normals estimated by a jointly trained NeuS model, which notably increases the training time. SuGaR [10] encourages Gaussians to align with the surface and subsequently treats the Gaussians as a point cloud for Poisson Surface Reconstruction. While these methods successfully extract a mesh, they require an additional step of training a new set of Gaussians bound to the mesh faces. This two-stage process brings inconvenience to existing downstream applications, including GS-Avatars [20], GS-VR [12], and physics simulations [8, 6]. This limitation motivates us to design a more effective information pathway to supervise the learning of scene representations.
@@
Figure 1: The Gaussians and surface did not align well in the original 3DGS, whereas our hybrid representation explicitly restricts the Gaussians to the mesh faces. Our method benefits from the high-quality rendering of 3DGS and the controllability of a mesh.
In this paper, we propose a novel approach that directly learns mesh and Gaussians from images in an end-to-end manner. Specifically, we utilize a differentiable marching module to extract a mesh, then derive Gaussians from the marched faces for strict alignment and employ the 3DGS workflow for rendering. We make the Gaussian covariances adaptive to irregular faces and use a neural network to robustly learn Gaussian colors as the mesh is jointly optimized. This resultant scene representation benefits from the rendering quality and efficiency of Gaussians, as well as the controllability of the mesh. Furthermore, the time required to extract a scene mesh is reduced to less than 2 hours on a single GPU, compared to other methods that rely on Neural SDFs and work on neural radiance fields [15, 27], which require more than one day of training.
@@
  • We enhance the existing differentiable mesh extraction method to broaden its applicability to outdoor scenes.
  • We introduce a scheme that adaptively adjusts the covariances of Gaussians based on their associated mesh faces and learns the colors of Gaussians through a neural network, thereby enhancing the robustness of the optimization process.
Experimental results demonstrate that our method can efficiently extract a scene representation with explicit geometry and achieve high-quality rendering. Furthermore, we have showcased its application in scene modification and adaptation to scene updates.
2 Related Work
2.1 Scene Reconstruction via Radiance Field
@@
2.2 Image-based Surface Reconstruction
@@
Several studies have achieved mesh extraction using 3DGS models. DreamGaussian [25] evaluates the density value of grid nodes using a mixture of neighborhood Gaussians and employs the Marching Cubes algorithm [16] to extract the isosurface. However, similar to NeRFs, the resulting density field may not accurately reflect the surface, leading to unsatisfactory surface quality in complex scenes. NeuSG [3] jointly learns NeuS and 3DGS models while introducing a regularization term to ensure consistency between the Gaussian orientation and the surface normal predicted by NeuS. This joint optimization significantly extends the training duration to over ten hours. Most recently, SuGaR [10] regularizes Gaussians using derived SDF values to align them with the surface and then treats the Gaussians as a point cloud for Poisson Surface Reconstruction. However, SuGaR requires the training of an additional set of Gaussians that are bound to the extracted mesh faces. In contrast, our approach simultaneously learns a hybrid representation of Gaussians and mesh from images.
3 Preliminary
3D Gaussian Splatting (3DGS) [14] uses a mixture of 3D Gaussians to model scene visuals. Each component is a Gaussian measure in 3D space, represented as $\exp\left\{-\frac{1}{2}(\boldsymbol{x}-\boldsymbol{\mu})^{T}\boldsymbol{\Sigma}^{-1}(\boldsymbol{x}-\boldsymbol{\mu})\right\}$, where $\boldsymbol{\mu}$ is the center position (mean) and $\boldsymbol{\Sigma}$ is the 3D covariance matrix. The visual information is characterized by opacity $\alpha$ and color $\boldsymbol{c}$, represented by spherical harmonics (SHs).
Physical properties that affect how a ray of light passes through the density field can be computed via integration through the density function. This naturally leads to an imaging model that can efficiently compute images based on Gaussian components by analytically solving the integration via projection:
@@ 3 Preliminary
where $\boldsymbol{\Sigma}^{\prime}$ is the projected 2D covariance matrix, $\boldsymbol{W}$ represents the current viewing transformation, and $\boldsymbol{J}$ is the Jacobian for an affine approximation to the projective transformation [32]. Additionally, the projected Gaussian center $\boldsymbol{\mu}_{i}^{\prime}\in\mathbb{R}^{2}$ results from the standard projective transformation.
To query the color of a pixel of interest at location $\boldsymbol{x}^{\prime}\in\mathbb{R}^{2}$, the sorted list of contributing Gaussians, $\mathcal{N}$, is determined based on their depths. The color of this pixel is then computed as follows:
$\boldsymbol{C}=\sum_{i\in\mathcal{N}}\boldsymbol{c}_{i}\alpha_{i}^{\prime}\prod_{j=1}^{i-1}(1-\alpha_{j}^{\prime}),$ (2)
@@
$\alpha_{i}^{\prime}=\alpha_{i}\times\exp\left\{-\frac{1}{2}(\boldsymbol{x}^{\prime}-\boldsymbol{\mu}_{i}^{\prime})^{T}\boldsymbol{\Sigma}_{i}^{\prime-1}(\boldsymbol{x}^{\prime}-\boldsymbol{\mu}_{i}^{\prime})\right\},$ (3)
@@
Using images from known camera poses, the parameters of the Gaussian components are adjusted to align rendered and observed images. The optimization variables for covariance matrix $\boldsymbol{\Sigma}$ include a quaternion $\boldsymbol{q}$ and a 3D vector $\boldsymbol{s}$, which relate to the objective through a change of variables:
@@ 4 Method
The proposed method aims to obtain set of Gaussians that strictly align with object surfaces. These surfaces are constructed with triangle meshes, and the Gaussians are precisely defined in relation to the mesh faces. This method includes modular differentiable steps, allowing the model to learn from image observations as in original 3DGS.
An overview of our method is illustrated in Fig. 2. The components of the proposed learnable 3D scene model consist of: (1) a learnable SDF grid, whose nodes represent the signed distances to the surfaces in the scene; (2) a mesh derived from the SDF grid; (3) Gaussians bound to the simplices (triangles in this work) of the mesh; (4) an appearance field represented by a neural network.
In the model workflow, the relationships between these components are as follows: The mesh (2) is derived from the SDF grid (1) using a differentiable marching algorithm. The Gaussians (3) are then created and populated with queries to the network (4) for rendering an image, which is used to compute the photometric loss.
4.1 Differentiable Surface Extraction
The learnable Signed Distance Function (SDF) $s:\mathbb{R}^{3}\mapsto\mathbb{R}$ is parameterized by scalar values at grid nodes $G=\{\boldsymbol{n}_{i},\dots,\boldsymbol{n}_{N}\}$. We adopt the recently proposed FlexiCubes [23] as the differentiable surface extractor, which improves the original Marching Cubes algorithm by introducing learnable per-node deformations and weights for cell interpolation and quadrilateral splitting. This process yields a triangle mesh, characterized by vertices $\mathcal{V}=\{\boldsymbol{v}_{1},\dots,\boldsymbol{v}_{V}\}\subset\mathbb{R}^{3}$ and faces =$\mathcal{F}=\{\boldsymbol{f}_{1},\dots,\boldsymbol{f}_{F}\}$, where each $\boldsymbol{f}$ denotes a triplet of indices within $[1,\dots,V]$.
Existing differentiable marching methods are primarily used for single objects. A critical factor behind this limitation is the exponential increase in storage and computational demands required for processing all cells as the scene scope widens. To handle real-world indoor scenes and even outdoor scenes, we use three techniques as follows:
Foreground-Background Grids.
We adopt two SDF grids to cover the pre-defined foreground and background domains, $D_{fg},D_{bg}\subset\mathbb{R}^{3}$, respectively. Cells in the background grid located within $D_{fg}$ are skipped during the marching procedure. For each training step, the resulting mesh is a composite of two meshes, each marched from one of the two grids.
Visible Node Optimization.
@@ 4.1 Differentiable Surface Extraction
As the camera only captures appearances within its current viewing frustum, it is unnecessary to extract mesh faces for the entire scene at each training iteration. Therefore, our focus shifts exclusively to the visible grid nodes, $\hat{G}$, calculated by:
$\hat{G}=\left\{\boldsymbol{n}\in G\big{|}\text{NDC}(\boldsymbol{n})\in[-1,1]^{3}\right\}$ (5)
where $\text{NDC}(\boldsymbol{n})$ is the transformation from world-space coordinates $\boldsymbol{n}$ to their normalized device coordinates (NDC) via the projection matrix. The marching algorithm is only applied to cells whose eight corner nodes are within $\hat{G}$.
Coarse-to-fine Grid.
Early use of a high-resolution grid in training often leads to numerous redundant faces, causing unstable learning and out-of-memory issues. Therefore, we start with a lower-resolution grid and increase the resolution several times (typically four times) during training until reaching the desired resolution. The enhancement of grid resolution is achieved through linear interpolation between the nodes of the existing grid and those of the refined grid. More details can be found in the appendix.
@@
Figure 3: Sub-figure (a) defines a local coordinate frame, while (b) illustrates applying linear transformation $\boldsymbol{M}$ to let Gaussians adapt to the irregular triangle. See Sec. 4.2 for more details.
4.2 Explicitly Surface-aligned and Adaptive Gaussians
As shown in Fig. 3, we associate each triangular face with $K$ Gaussians for a given extracted mesh $(\mathcal{V},\mathcal{F})$. For a single face defined by vertices $\{\boldsymbol{v}_{1},\boldsymbol{v}_{2},\boldsymbol{v}_{3}\}$, the centers of the Gaussians are computed as follows:
$\boldsymbol{\mu}_{i}=\boldsymbol{bc}_{i}[\boldsymbol{v}_{1},\boldsymbol{v}_{2},\boldsymbol{v}_{3}]^{T},\quad i=1,2,\dots,K,$ (6)
where $\boldsymbol{\mu}_{i}$ represents the center of the $i$-th Gaussian and $\boldsymbol{bc}_{i}$ denotes the predefined barycentric coordinates. The configurations of $\boldsymbol{bc}$ for different values of $K$ are detailed in the appendix.

### v2

@@ Abstract
Accurately reconstructing a 3D scene including explicit geometry information is both attractive and challenging. Geometry reconstruction can benefit from incorporating differentiable appearance models, such as Neural Radiance Fields and 3D Gaussian Splatting (3DGS). However, existing methods encounter efficiency issues due to indirect geometry learning and the paradigm of separately modeling geometry and surface appearance. In this work, we propose a learnable scene model that incorporates 3DGS with an explicit geometry representation, namely a mesh. Our model learns the mesh and appearance in an end-to-end manner, where we bind 3D Gaussians to the mesh faces and perform differentiable rendering of 3DGS to obtain photometric supervision. The model creates an effective information pathway to supervise the learning of both 3DGS and mesh. Experimental results demonstrate that the learned scene model not only achieves state-of-the-art efficiency and rendering quality but also supports manipulation using the explicit mesh. In addition, our model has a unique advantage in adapting to scene updates, thanks to the end-to-end learning of both mesh and appearance.
Index Terms:
@@
3D reconstruction, Gaussian splatting, Hybrid representation, Mesh-guided manipulation
I Introduction
@@
Mesh is often the preferred choice in numerous geometry processing scenarios, as it carry explicit surface information. When building a 3D scene model that includes a mesh, it is challenging to directly learn from visual observation due to the complicated pipeline required to produce the synthetic views. Recently, neural surface reconstruction methods [28, 32] have been developed to represent scene geometry as an implicit signed distance function (SDF). These methods establish a connection with Neural Radiance Fields (NeRFs) [17], enabling effective image-based supervision for geometry learning. However, the geometry is learned indirectly, as it involves converting the SDF to a radiance field. Additionally, NeRF-based learning requires expensive ray-marching in the volumetric rendering. These factors lead to efficiency challenges in learning geometry and appearance.
@@
3D Gaussian Splatting (3DGS) [14] has recently gained popularity due to its ability to render photorealistic images significantly faster than NeRFs. 3DGS employs an explicit representation called Gaussians, which represent anisotropic ellipsoids in 3D space. Gaussians enable efficient rendering due to their explicit nature and the highly parallelized splatting procedure. Despite the advantages in rendering quality and speed, it is not straightforward to reconstruct object surfaces in the traditional form of Gaussians. This is because the 3DGS model imposes no explicit constraints on the structure of the Gaussians. As illustrated in Fig. 1, Gaussians that result in visually appealing rendering may be geometrically inaccurate – they can be inconsistent with any 2D manifold. Recently, some methods have incorporated regularization into the learning process of 3DGS. NeuSG [3] refines Gaussians using surface normals estimated by a jointly trained NeuS model [28], which notably increases the training time. SuGaR [10] encourages Gaussians to align with the surface and subsequently converts them into a point cloud for Poisson Surface Reconstruction. Once a coarse mesh is extracted, an additional stage is required to learn surface appearance from scratch using a new set of Gaussians. This two-stage pipeline may introduce efficiency problems. It is worth noting that similar issues are encountered in existing downstream applications [20, 12, 8, 6], which perform meshing and Gaussian training separately. These challenges motivate us to design a more effective information pathway to supervise the learning of scene representations.
@@
Figure 1: The Gaussians and the underlying surface did not align well in the original 3DGS [14], whereas our hybrid representation explicitly restricts the Gaussians to the mesh faces. Our method benefits from the high-quality rendering of 3DGS and the controllability of a mesh.
In this paper, we present a technique to explicitly bind Gaussians to mesh faces, enabling their simultaneous learning. Specifically, the mesh is extracted from a learnable SDF of objects using differentiable marching techniques. The 3D Gaussians are then constructed as disks that are regularly attached to the mesh faces, as shown in Fig. 1 (b). Since the 3DGS model is bound to the zero-level set of the SDF representing object surfaces, the photometric supervision applied to the 3DGS model simultaneously guides the reconstruction of both the appearance and geometry of a scene. In addition, we model the background elements using the original 3DGS, which is jointly learned with our proposed surface-bound Gaussians. This approach makes our method more practical compared to existing methods that require inconvenient foreground masks [19, 23]. As the mesh faces vary during training, the Gaussians dynamically bound to them struggle to learn per-Gaussian colors as effectively as in the original 3DGS. To address this issue, we employ a neural network to learn the scene’s appearance and predict Gaussian colors for rendering.
Extensive experiments show that our method achieves state-of-the-art rendering accuracy and produces high-quality meshes in indoor and outdoor scenes. Furthermore, it reduces total reconstruction time by 50% compared to recent methods and is over 10x faster than methods that rely on Neural SDFs and NeRFs [15, 28]. We also demonstrate that our method has a unique advantage in adapting to scene updates compared to other two-stage methods [10, 27].
@@
  • We propose a technique to model background appearance, extending the applicability of differentiable mesh extraction methods to outdoor scenes without the need for foreground masks.
  • We introduce neural appearance model that predicts Gaussians colors, enhancing the robustness of the optimization process.
II Related Work
II-A Scene Reconstruction via Radiance Field
@@
II-B Image-based Surface Reconstruction
@@
Several studies have achieved mesh reconstruction using 3DGS models. DreamGaussian [25] evaluates the density value of grid nodes using a mixture of neighborhood Gaussians and employs the Marching Cubes algorithm [16] to extract the isosurface. However, similar to NeRFs, the resulting density field may not accurately reflect the surface, leading to unsatisfactory surface quality in complex scenes. NeuSG [3] jointly learns NeuS and 3DGS models while introducing a regularization term to ensure consistency between the Gaussian orientation and the surface normal predicted by NeuS. This joint optimization significantly extends the training duration to over ten hours. Most recently, SuGaR [10] regularizes Gaussians using derived SDF values to align them with the surface and then treats the Gaussians as a point cloud for Poisson Surface Reconstruction. However, SuGaR requires the training of an additional set of Gaussians that are bound to the extracted mesh faces.
We also highlight the differences between our method and the recent works SuGaR [10] and GaMeS [27], which both construct a hybrid representation of Gaussians and mesh. Existing works first extract the mesh or directly use the ground-truth mesh and then learn Gaussian-based appearance from scratch. In contrast, our approach simultaneously learns such a hybrid representation in an end-to-end manner. We will demonstrate that our method achieves a more efficient pipeline and better performance.
III Preliminary
III-A 3D Gaussian Splatting
3D Gaussian Splatting (3DGS) [14] represents the 3D scene using a collection of Gaussians, each defined as a density function in $\mathbb{R}^{3}$:
$G(\boldsymbol{x})=\exp\left\{-\frac{1}{2}(\boldsymbol{x}-\boldsymbol{\mu})^{T}\boldsymbol{\Sigma}^{-1}(\boldsymbol{x}-\boldsymbol{\mu})\right\}$ (1)
where $\boldsymbol{\mu}$ is a center position (mean), and $\boldsymbol{\Sigma}$ denotes a 3D covariance matrix. To model scene appearance, each Gaussian is associated with an opacity $\alpha\in[0,1]$, and a view-dependent color $\boldsymbol{c}$, which is parameterized with a set of Spherical Harmonics coefficients (SHs) $\{\boldsymbol{y}_{l}^{m}\in\mathbb{R}^{3}\}_{0\leq l\leq l_{\text{max}}}^{-l\leq m\leq l}$ [7].
The interaction of a ray with a Gaussian can be characterized by integrating the Gaussian’s density function $G(\boldsymbol{x})$. This leads to an efficient imaging model that projects Gaussians into 2D image space, resulting in 2D Gaussians with covariance calculated as:
@@ III Preliminary
where $\boldsymbol{W}$ represents the current viewing transformation, and $\boldsymbol{J}$ is the Jacobian for an affine approximation to the projective transformation [33].
To query the color of a pixel at location $\boldsymbol{x}^{\prime}\in\mathbb{R}^{2}$ on image $\mathbf{I}$, a sorted list $\mathcal{S}$ of contributing Gaussians is determined based on their depths. The color of this pixel is then computed as follows:
$\mathbf{I}(\boldsymbol{x}^{\prime})=\sum_{i\in\mathcal{S}}\boldsymbol{c}_{i}\alpha_{i}^{\prime}\prod_{j=1}^{i-1}(1-\alpha_{j}^{\prime}),$ (3)
@@
$\alpha_{i}^{\prime}=\alpha_{i}\exp\left\{-\frac{1}{2}(\boldsymbol{x}^{\prime}-\boldsymbol{\mu}_{i}^{\prime})^{T}\boldsymbol{\Sigma}_{i}^{\prime-1}(\boldsymbol{x}^{\prime}-\boldsymbol{\mu}_{i}^{\prime})\right\},$ (4)
@@
Using images from known camera poses, the parameters of the Gaussians are adjusted to align rendered and observed images $\mathbf{I}_{\text{GT}}$ by minimizing the photometric loss:
$\underset{\alpha,\boldsymbol{\mu},\boldsymbol{c},\boldsymbol{\Sigma}}{\text{minimize }}\mathcal{L}_{\text{photo}}(\mathbf{I},\mathbf{I}_{\text{GT}}).$ (5)
The optimization variables for covariance matrix $\boldsymbol{\Sigma}$ include a quaternion $\boldsymbol{q}$ and a 3D vector $\boldsymbol{s}$, which relate to the objective through a change of variables:
@@
III-B Signed Distance Function
The Signed Distance Function (SDF) is a scalar field that defines the distance from any point in space to the nearest surface of an object. The sign of the distance indicates whether the point is inside or outside the object. Formally, given a point $\boldsymbol{x}$ in 3D space, the SDF value is defined as:
$D(\boldsymbol{x})=\begin{cases}+d(\boldsymbol{x},\partial\Omega)&\text{if }\boldsymbol{x}\in\mathbb{R}^{3}\setminus\Omega\\ -d(\boldsymbol{x},\partial\Omega)&\text{if }\boldsymbol{x}\in\Omega\end{cases}$ (7)
where $d(\boldsymbol{x},\partial\Omega)$ is the Euclidean distance from $\boldsymbol{x}$ to the surface $\partial\Omega$ of the object $\Omega$.
Directly learning a mesh from images is challenging. A common paradigm is to optimize an parameterized SDF and then extract a triangle mesh approximating the level set of that function. Notable examples include recent works that use differentiable marching techniques [30, 23, 19] and the conversion from SDF to neural radiance fields [15, 28, 29].
IV Method
@@
The proposed method aims to obtain a set of Gaussians that strictly align with object surfaces. Specifically, these surfaces are constructed with triangle meshes, and the Gaussians are precisely defined in relation to the mesh faces.
As illustrated in Fig. 2, our method includes modular differentiable steps, allowing the model to learn from images as in the original 3DGS. Let the 3D scene of interest be within a region $U\subset\mathbb{R}^{3}$. The model consists of:
  1. 1.
    a scalar field of signed distance function $D:U\mapsto\mathbb{R}$, which relates locations to the presence of objects;
  2. 2.
    a triangle mesh $G=(\mathcal{V},\mathcal{F})$ derived from $D$, where vertices $\mathcal{V}=\{\boldsymbol{v}_{1},\dots,\boldsymbol{v}_{V}\}\subset\mathbb{R}^{3}$ and faces =$\mathcal{F}=\{\boldsymbol{f}_{1},\dots,\boldsymbol{f}_{F}\}$, with each $\boldsymbol{f}$ denoting a triplet of indices within $[1,\dots,V]$;
  3. 3.
    a constrained 3DGS model where the Gaussians are attached to the faces in $\mathcal{F}$;
  4. 4.
    a neural appearance model, which maps a Gaussian center to spherical harmonics coefficients.
IV-A Signed Distance Function and Differentiable Mesh Computation
The geometry of the scene consists of object surfaces, represented as the zero-level sets of the signed distance function, denoted by $\{\boldsymbol{x}|D(\boldsymbol{x})=0\}$. We represent $D$ using an explicit format: for a set of points $\{\boldsymbol{x}_{n}\}_{n=1}^{N}$, which correspond to the nodes of a regular 3D grid, the function values $\{s_{n}\}_{n=1}^{N}$ are specified. For a generic location $\boldsymbol{x}$, $D(\boldsymbol{x})$ is computed by interpolation from the eight nodes of the grid cell enclosing $\boldsymbol{x}$. Therefore, the value set $\{s_{n}\}_{n=1}^{N}$ forms a parametric representation of $D$.
Given the SDF, the object surfaces are implicitly defined by its zero-level set. However, a practical 3D scene model requires explicit geometry to be incorporated into the computational pipelines for rendering and learning. A traditional method to recover a discrete surface representation for a given SDF is the Marching Cubes algorithm [16], which results in a triangle mesh consistent with a 2D manifold. In contrast, this work involves not just recovering a surface but also learning the SDF. This requires the model to adjust the SDF parameters $\{s_{n}\}_{n=1}^{N}$ to make the resultant surfaces vary according to desired updates. Therefore, we adopt the recently proposed FlexiCubes [23] as the differentiable surface extractor.
Additionally, we address two substantial challenges of modeling outdoor and realistic scenes when learning the SDF: (1) For real-world scenes, especially unbounded outdoor environments, the scale of the background ($\mathbb{R}^{3}\setminus U$) is significantly large and costly to represent using an explicit mesh unified with the foreground $U$. Existing methods often manually exclude the background using known foreground segmentation masks [19, 23]. This approach is inconvenient in practice, and visual modeling of the background is still necessary for tasks such as scene image synthesis. (2) As the scale of $U$ increase, the memory required to execute differentiable marching algorithms grows significantly.
We address challenge (1) by incorporating an additional set of Gaussians to handle the visual components from the background. Specifically, we construct background Gaussians $\mathcal{GS}_{\text{bg}}=\{\alpha_{k},\boldsymbol{\mu}_{k},\boldsymbol{c}_{k},\boldsymbol{\Sigma}_{k}\}$ using the original form of 3DGS. Due to their explicit nature, $\mathcal{GS}_{\text{bg}}$ can be directly combined with the foreground Gaussians $\mathcal{GS}_{\text{fg}}$ attached to the mesh (detailed in the next section), enabling rendering and training.
For challenge (2), we apply two schemes in the use of the SDF grid to mitigate the storage and computational demands, thus enabling a larger scale of $U$:
Visible Node Optimization
@@ IV-A Signed Distance Function and Differentiable Mesh Computation
During each training iteration, only the scene elements within a single camera’s viewing frustum are learning. Therefore, we extract mesh faces only from the visible grid nodes in $\{\boldsymbol{x}_{n}\}_{n=1}^{N}$ using the camera’s intrinsic and extrinsic parameters.
Coarse-to-fine Grid
Starting with a high-resolution grid can lead to redundant faces, causing unstable learning and memory issues. To address this, we begin with a low-resolution grid and progressively increase the resolution during training. The resolution is enhanced by linearly interpolating between the nodes of the current grid and the refined grid.
More technical details can be found in the appendix.
@@
Figure 3: Sub-figure (a) shows using $K=1,3,6$ Gaussians to represent the appearance of a triangle face, (b) defines a local coordinate frame, and (c) illustrates applying linear transformation $\boldsymbol{M}$ to let Gaussians adapt to the irregular triangle. See Sec. IV-B for more details.
IV-B Gaussians Consistent with Scene Geometry
As shown in Fig. 3 (a), we employ $K$ Gaussians to represent the surface appearance within a triangle $(\boldsymbol{v}_{1},\boldsymbol{v}_{2},\boldsymbol{v}_{3})$, where $K$ depends on desirable level of detail.
The centers of the Gaussians are placed at specific barycentric coordinates within the triangle. Let these barycentric coordinates be $\boldsymbol{\xi}_{1},\boldsymbol{\ldots},\boldsymbol{\xi}_{K}\in\mathbf{S}^{2}$, where $\mathbf{S}^{2}$ is the 2-simplex: $\mathbf{S}^{2}=\{(x_{1},x_{2},x_{3})|x_{n}\geq 0,x_{1}+x_{2}+x_{3}=1\}$. In the world coordinate system, the center of a Gaussian is computed as:
$\boldsymbol{\mu}_{k}=[\boldsymbol{v}_{1},\boldsymbol{v}_{2},\boldsymbol{v}_{3}]\cdot\boldsymbol{\xi}_{k}.$ (8)
Note that although the barycentric coordinates $\{\boldsymbol{\xi}_{k}\}$ are fixed, the centers $\{\boldsymbol{\mu}_{k}\}$ depend on the mesh vertices $\boldsymbol{v}_{1},\boldsymbol{v}_{2},\boldsymbol{v}_{3}$, which in turn depend on the SDF field. The configurations of $\{\boldsymbol{\xi}_{k}\}$ for different $K$ values can be found in the appendix.